Suppose a particular consumer has 8 birr to be spent on two goods A and B the unit price of goods A is 2 birr and the unit price of B is 1 birr.
The marginal utility (MU) she gets from consumption of the goods is given below
1
Expert's answer
2021-09-05T17:54:21-0400

a. The utility is maximized when "\\frac{MUA}{PA}=\\frac{MUB}{PB}\n\n\u200b"

 and

"PA \times QA + PB \times QB = 8."

Both equities are true if QA = 2 units and QB = 4 units.


b. The total utility at the utility maximization level is:

"TU = (36 + 24) + (30 + 22 + 16 + 12) = 140 units."